Circular No. 6615

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) R. Meier, University of Hawaii; H. Matthews, Joint Astronomy Centre

(JAC), Hilo, and National Research Council of Canada; T. Owen and D. Jewitt, University of Hawaii; D. Gautier, N. Biver, D. Bockelée-Morvan, and J. Crovisier, Observatoire de Paris, Meudon, report the first detection of HDO in comet C/1995 O1 (see IAUC 6604) with the James Clerk Maxwell Telescope at Mauna Kea on Apr. 4.8 UT: “The integrated line intensity of the 1_{01}–0_{00} HDO line at 464.925 GHz was (0.64 ± 0.11) K km/s on the antenna temperature scale. We also detected two weak lines of methanol at 464.835 GHz (9_{2}–9_{1} transition, T_{A} = 1.21 ± 0.11 K km/s) and 457.000 MHz (11_{2}–11_{1} transition, T_{A} = 2.37 ± 0.11 K km/s). Assuming a rotational temperature of 80 K, we derive a preliminary HDO production rate of 4 × 10^{26} molecules/s. This makes the HDO/H_{2}O ratio ∼ 2 × 10^{−4}, assuming a water-production rate Q(H_{2}O) = 2 × 10^{30} molecules/s (IAUC 6589). The corresponding D/H ∼ 10^{−4} is consistent with other values for hydrogen in condensed matter in the outer solar system and in seawater on earth, and is well above the values reported for H_{2} in the interstellar medium or in outer-planet atmospheres.”

SUPERNOVA 1997bo IN ANONYMOUS GALAXY W.-d. Li, Beijing Astronomical Observatory (BAO), reports that the

field of SN 1997bo (cf. IAUC 6612) was observed numerous times in the past two months by the BAO Supernova Survey with the BAO 0.6-m reflector (+ unfiltered CCD), yielding the following estimated magnitudes: Feb. 10, [17.5; 15, [18.0; Mar. 1, 15.7 ± 0.2; 5, 15.8 ± 0.2; 15,15.9 ± 0.2; 27, 16.0 ± 0.2; Apr. 3, 16.1 ± 0.2; 4, 16.1 ± 0.2. The host galaxy is faint and irregular, making it difficult to determine the position of its nucleus; there is a relatively bright H II region east of its nucleus, close to SN 1997bo. A precise measurement of the supernova from an Apr. 4 CCD frame yields α = 12^{h}19^{m}52^{s}.66, δ = +7^{o}43′49′′.7 (equinox 2000.0), which is 2′′.4 east and 0′′.7 north of the galaxy’s center.

COMET 46P/WIRTANEN Visual m_{1} magnitude estimates: Mar. 6.80 UT, 9.6 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); 31.84, 10.0 (Bouma); Apr. 2.79, 10.0 (K. Sarneczky, Raktanya, Hungary, 0.44-m reflector); 6.84, 9.6 (A. Kammerer, Ettlingen, Germany, 0.20-m reflector).

1997 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green